OpenStudy (twizttiez):

The system of equations is coincident. What are the missing values? -2x+5y=-12 -6x+___y=___?

OpenStudy (twizttiez):

@Algorithmic

OpenStudy (twizttiez):

@satellite73

OpenStudy (twizttiez):

@triciaal

OpenStudy (triciaal):

coincident i think means lie on the same line one is a factor of the other

OpenStudy (anonymous):

\[-2\times \text{what number}=-6\]?

OpenStudy (triciaal):



OpenStudy (twizttiez):

3

OpenStudy (twizttiez):

Is?

OpenStudy (triciaal):

so the 2nd is 3* first

OpenStudy (twizttiez):

??

OpenStudy (twizttiez):

@Loser66

OpenStudy (twizttiez):

So is my answer -6x+(-2)y=3?

OpenStudy (triciaal):

-2x+5y=-12 eqn 1 -6x+___y=___? eqn 2 the eqn 2 is 3 times the eqn 1 3(-2x + 5y) = 3(-12)

OpenStudy (twizttiez):

Does that mean i was right or wrong?

OpenStudy (twizttiez):

@satellite73

OpenStudy (twizttiez):

@Algorithmic

OpenStudy (twizttiez):

@triciaal

OpenStudy (twizttiez):

@sleepyjess

OpenStudy (triciaal):

3(-2x + 5y) = 3(-12) what do you get when you expand? rephrased get rid of the parenthesis what do you have ?